OpenStudy (anonymous):

h=-16t^2+116t+30 how do i find maximum height and how long it took to get there?

OpenStudy (anonymous):

Are you familiar with completing the square or finding derivatives? @revernt

OpenStudy (anonymous):

@revernt Are you there?

OpenStudy (anonymous):

I'm pretty sure there's a formula for it? x=-b/2a, or in this case, t=-116/2(-16) is the highest point, simplify: -116/-32=3.625 (that would be the time). Substitute it into the equation: h(3.625)=-16(3.625)^2+116(3.625)+30=-210.25+420.5+30=240.25 (the height).

OpenStudy (anonymous):

@revernt

OpenStudy (anonymous):

thanks! @mangnagelytis

OpenStudy (anonymous):

You're welcome.